# Radial and Non-Radial Solution Structures for Quasilinear Hamilton–Jacobi–Bellman Equations in Bounded Settings

### v2

@@
Abstract
We study quasilinear Hamilton–Jacobi–Bellman equations on bounded smooth convex domains. We show that the quasilinear Hamilton–Jacobi–Bellman equations arise naturally from stochastic optimal control problems with exit-time costs. The PDE is obtained via dynamic programming applied to controlled Itô diffusions, providing both a probabilistic interpretation and a rigorous derivation. Our result establishes existence and uniqueness of positive classical solutions under sub-quadratic growth conditions on the source term. The constructive proofs, based on monotone iteration and barrier techniques, also provide a framework for algorithmic implementation with applications in production planning and image restoration. We provide complete detailed proofs with rigorous estimates and establish the connection to stochastic control theory through the dynamic programming principle.
Mathematics Subject Classification (2020): 35J60, 35J70, 35B40, 49L25, 65N12.
Keywords and phrases: Hamilton–Jacobi–Bellman equations; quasilinear elliptic PDEs; nonradial solutions; monotone iteration; production planning; image restoration.
1 Introduction
In this paper we investigate the existence, uniqueness, and regularity of solutions to the quasilinear Hamilton–Jacobi–Bellman (HJB) equation
@@ 1 Introduction
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded $C^{2}$ convex domain with smooth boundary $\partial\Omega$, $\sigma>0$ represents the diffusion coefficient, $\alpha\in(1,2]$ is the cost exponent, and $g\geq 0$ is a constant,
$p:=\frac{\alpha}{\alpha-1}\in[2,\infty),\qquad C_{\alpha}:=\frac{\alpha-1}{\alpha^{\frac{\alpha}{\alpha-1}}}>0.$
@@
The parameter $p$ is conjugate to $\alpha$ via the relation $1/\alpha+1/p=1$, and the constant $C_{\alpha}$ arises from the Legendre transform of the control cost functional.
The problem (1.1) was first introduced by Lasry and Lions [4] as a mathematical model, and later studied from the perspective of applications in production planning in our work [1] (see also [3]).
The motivation for considering (1.1) is twofold. First, the equation arises in image restoration problems, as discussed in our recent work [2], where the parameter $\alpha$ close to $1$ is used for contrast of a image (as a constraint) in the image model. Second, from a purely mathematical perspective, the case $\alpha\in(1,2)$ remains largely unexplored for general bounded domains $\Omega$ (see [2, 3]).
Our main result establishes the existence and uniqueness of positive solutions to the Dirichlet problem (1.1) under sub-quadratic growth conditions on $h$.
Theorem 1.1.
@@
Let $h:\overline{\Omega}\rightarrow[0,\infty)$ be continuous and sub-quadratic (i.e., there exist constants $C_{h}>0$ and $p:=\frac{\alpha}{\alpha-1}\in[2,\infty)$ with $\alpha\in(1,2]$ such that $h(y)\leq C_{h}(1+|y|^{p})$ for all $y\in\overline{\Omega}$), and let $g\in[0,\infty)$ be a constant boundary datum. Then the Dirichlet problem (1.1) admits a unique positive solution $V\in C^{2}(\Omega)\cap C(\overline{\Omega})$. Moreover, if $g>0$ on $\partial\Omega$ and $h>0$ in $\Omega$, then $V>0$ in $\overline{\Omega}$.
Remark 1.2.
The regularity assumption $h\in C(\overline{\Omega})$ can be relaxed to $h\in L^{\infty}(\Omega)$ with appropriate modifications to the proof.
Finally, we emphasize that in order to apply these theoretical results in practice (for instance, in production planning or image restoration), one requires proofs that can be implemented algorithmically. A natural approach is to employ successive approximation techniques combined with the construction of sub- and super-solutions, as successfully implemented in our previous works [2, 3].
2 Proof of Theorem 1.1
The proof proceeds in four steps: (1) construction of ordered sub- and super-solutions, (2) comparison principle, (3) existence via monotone iteration, and (4) uniqueness. Each step is presented with complete technical details.
@@
2.1 Step 1: Construction of ordered positive sub- and super-solutions
We construct a positive sub-solution $V_{-}$ and a positive super-solution $V_{+}$ such that
$V_{-}\leq V_{+}\ \text{in}\ \Omega,\qquad V_{-}\leq g\leq V_{+}\ \text{on}\ \partial\Omega.$
Since $g$ is constant, the construction simplifies significantly.
@@ 2.1 Step 1: Construction of ordered positive sub- and super-solutions
Construction based on the torsion function. Let $\phi\in C^{\infty}(\Omega)\cap C(\overline{\Omega})$ be the unique solution to the Poisson (torsion) problem
@@
$\Delta\phi=1\quad\text{in }\Omega,\qquad\phi=0\quad\text{on }\partial\Omega.$ (2.1)
Since $\Omega$ is a bounded $C^{2}$ domain, classical elliptic regularity theory guarantees that $\phi\in C^{\infty}(\Omega)\cap C(\overline{\Omega})$. Moreover, by the strong maximum principle, $\phi>0$ in $\Omega$, and by the Hopf lemma, $\frac{\partial\phi}{\partial\nu}<0$ on $\partial\Omega$, where $\nu$ denotes the outward unit normal.
Sub-solution. Let
$V_{-}(y):=g,\quad y\in\overline{\Omega}.$
Then $\nabla V_{-}\equiv 0$ and $\Delta V_{-}\equiv 0$, so the differential operator evaluated at $V_{-}$ gives
@@
because $h\geq 0$. Hence $V_{-}$ is a (classical) sub-solution in $\Omega$, and $V_{-}=g$ on $\partial\Omega$. Since $g\geq 0$, $V_{-}$ is positive.
Super-solution. Fix a constant $B>0$ to be determined, and define
$V_{+}(y):=g+B\,\phi(y),\qquad y\in\overline{\Omega}.$ (2.2)
By construction, $V_{+}=g$ on $\partial\Omega$ since $\phi=0$ there, and $V_{+}\geq g\geq 0$ in $\Omega$ because $\phi>0$ in $\Omega$. Computing the derivatives:
$\nabla V_{+}=B\,\nabla\phi,\qquad\Delta V_{+}=B\,\Delta\phi=-\,B,$
where we used (2.1). Substituting into the differential operator yields
$-\frac{\sigma^{2}}{2}\Delta V_{+}\;+\;C_{\alpha}\,|\nabla V_{+}|^{p}\;-\;h(y)$ $=\;\frac{\sigma^{2}}{2}\,B\;+\;C_{\alpha}\,B^{p}\,|\nabla\phi(y)|^{p}\;-\;h(y)$
$\geq\;\frac{\sigma^{2}}{2}\,B\;-\;h(y),$ (2.3)
where $p=\alpha/(\alpha-1)$ and $C_{\alpha}=(\alpha-1)/\alpha^{p}$, and the inequality holds because $|\nabla\phi|^{p}\geq 0$.
Since $\Omega$ is bounded and $h\in C(\overline{\Omega})$, we have
$H:=\max_{y\in\overline{\Omega}}h(y)<\infty.$
Therefore, from (2.3), if we choose
$B\;\geq\;\frac{2H}{\sigma^{2}},$ (2.4)
it follows that
@@
Thus $V_{+}$ is a classical super-solution to (1.1), satisfies $V_{+}=g$ on $\partial\Omega$, and $V_{+}>0$ in $\Omega$ (since $\phi>0$ in $\Omega$ and $g\geq 0$).
Order and positivity. By construction,
$V_{-}(y)=g\;\leq\;g+B\,\phi(y)=V_{+}(y)\quad\text{for all }y\in\Omega,$
with equality on $\partial\Omega$ since $\phi=0$ there. Both $V_{-}$ and $V_{+}$ are positive provided $g\geq 0$.
Conclusion. With
$V_{-}(y)\equiv g\quad\text{and}\quad V_{+}(y)=g+B\phi(y)\quad\text{for }B\geq 2H/\sigma^{2},$
we have constructed positive sub- and super-solutions satisfying $V_{-}\leq V_{+}$ in $\Omega$ and $V_{-}\leq g\leq V_{+}$ on $\partial\Omega$.
Remarks.
  • The sub-quadratic growth of $h$ ensures $H<\infty$ on the bounded domain $\Omega$, which is all we need to pick $B$ as above. No further regularity of $h$ is required.
  • If one prefers a radial barrier and $\Omega$ is $C^{2}$ and strictly contained in a ball $B_{R}(0)$, one may also take $V_{+}(y)=g+B\psi(y)$ with $\psi$ solving $\Delta\psi=1$ in $\Omega$, $\psi=0$ on $\partial\Omega$ (as above), which avoids geometric mismatch of a purely radial function with a general $\partial\Omega$.
2.2 Step 2: Comparison principle
The comparison principle is crucial for both the monotone iteration scheme and the uniqueness proof. We establish it through a perturbation argument combined with the maximum principle.
Lemma 2.1 (Comparison Principle).
Let $u,v\in C^{2}(\Omega)\cap C(\overline{\Omega})$ be, respectively, a subsolution and a supersolution to (1.1):
@@ 2.2 Step 2: Comparison principle
with $u\leq v$ on $\partial\Omega$. Then $u\leq v$ in $\Omega$.
Proof.
We argue by contradiction using a perturbation technique. Suppose that $\sup_{\Omega}(u-v)>0$. Since $u\leq v$ on $\partial\Omega$, the maximum of uvu-v must be attained at an interior point $y_{0}\in\Omega$, where
$(u-v)(y_{0})=\max_{\overline{\Omega}}(u-v)=:\delta>0,\qquad\nabla u(y_{0})=\nabla v(y_{0}),\qquad\Delta(u-v)(y_{0})\leq 0.$ (2.6)
Evaluating the inequalities (2.5) at $y_{0}$ and subtracting yields

### v3

@@
Abstract.
This paper establishes the existence, uniqueness, and global $C^{1,\beta}$ regularity of positive classical solutions to a class of quasilinear Hamilton–Jacobi–Bellman (HJB) equations with Dirichlet boundary conditions on bounded convex domains. The core technical contribution is a constructive existence proof based on a weighted linear monotone iteration scheme. This scheme’s stability and convergence are rigorously established through the construction of adaptive sub- and super-solutions leveraging the torsion function of the domain. Additionally, we provide a complete probabilistic derivation of the quasilinear PDE from the framework of controlled Itô diffusions, formally bridging the gap between stochastic optimal control theory and elliptic regularity analysis. Our results extend beyond the classical quadratic cost regime to the wider class of sub-quadratic growth source terms. Finally, we demonstrate the utility of this theoretical framework through high-precision numerical implementations in two distinct fields: stochastic production planning and nonlinear contrast enhancement in image restoration.
Key words and phrases:
Hamilton–Jacobi–Bellman equations, quasilinear elliptic PDEs, nonradial solutions, monotone iteration, production planning, image restoration
2020 Mathematics Subject Classification:
Primary 35J60; Secondary 35J70, 35B40, 49L25, 65N12
The final version of the article is published in journal (with volume, page numbers and DOI, when available).
1. Introduction
Nonlinear partial differential equations and optimal control techniques have played a central role in modern image processing and applied mathematics. Classical diffusion-based models such as the anisotropic diffusion of Perona and Malik [12] and the total variation framework of Rudin, Osher, and Fatemi [13] established the foundation for PDE-driven image enhancement and restoration. In parallel, the mathematical theory of Hamilton–Jacobi–Bellman (HJB) equations has been extensively developed in the context of stochastic control, with fundamental contributions by Lions [11], Lasry and Lions [10], Barles and Perthame [1], and the viscosity solution framework of Crandall, Ishii, and Lions [5]. The classical monographs of Gilbarg and Trudinger [7] and Fleming and Soner [6] provide comprehensive treatments of elliptic regularity and controlled diffusion processes, respectively.
In this paper we investigate the existence, uniqueness, and regularity of solutions to the quasilinear Hamilton–Jacobi–Bellman equation
@@ 1. Introduction
where $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded $C^{2}$ convex domain with smooth boundary $\partial\Omega$, $\sigma>0$ is the diffusion coefficient, $\alpha\in(1,2]$ is the cost exponent, and $g\geq 0$ is a constant boundary value. The parameters
$p:=\frac{\alpha}{\alpha-1}\in[2,\infty),\qquad C_{\alpha}:=\frac{\alpha-1}{\alpha^{\frac{\alpha}{\alpha-1}}}>0,$
@@
are linked through the conjugacy relation $1/\alpha+1/p=1$, and $C_{\alpha}$ arises naturally from the Legendre transform of the control cost functional.
The equation (1.1) was first introduced by Lasry and Lions [10] in the pioneering work on stochastic control with state constraints, where an additional discount factor is incorporated and the resulting solutions typically exhibit blow-up at the boundary. In contrast, our focus is on the Dirichlet problem on bounded convex domains, where the behavior of the solution and its gradient near the boundary is governed by the interplay between the diffusion coefficient $\sigma$ and the nonlinearity $C_{\alpha}|\nabla V|^{p}$.
More recently, this framework has found significant applications in production planning [2, 3] and variational image restoration [4]. The analysis is carried out on bounded domains when $\alpha=2$ [3] and in radially symmetric settings for all $\alpha\in\left(0,2\right]$ [4].
In particular, the regime $\alpha\in(1,2)$, which corresponds to a supra-quadratic gradient penalty $p>2$, generates a rich geometric structure that is highly effective for contrast enhancement.
This work’s central contribution is its unified treatment of broader problem settings than those addressed so far, supported by a mathematically rigorous and algorithmically constructive self-contained existence theory. By establishing the $C^{1,\beta}(\overline{\Omega})$ regularity for the case of sub-quadratic forcing $h$, we provide a missing link in the literature between the theoretical existence results of the 1980s and the high-performance numerical solvers required for modern industrial applications.
Significance and Novelty
Unlike standard treatments of the HJB equation that rely on the vanishing viscosity method or purely probabilistic arguments, our approach utilizes a weighted monotone linearization. This method not only proves existence but also yields a numerically stable iteration scheme that preserves the physical properties of the value function (such as positivity and concavity) at each step. This work establishes that for $\alpha\in(1,2]$, the geometric structure of the HJB solution in convex domains is robust enough to handle the supra-quadratic gradient penalty $p>2$, which is a regime classically avoided in simpler elliptic contexts.
The main results of this paper are established in the following two theorems, which characterize the structure of solutions to quasilinear Hamilton–Jacobi–Bellman equations in bounded settings.
Theorem 1.1 (Non-radial solutions).
@@
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded $C^{2}$ convex domain. Suppose $h:\overline{\Omega}\rightarrow[0,\infty)$ is a continuous function such that $\lim_{\left|x\right|\rightarrow\infty}\left(h\left(x\right)/\left|x\right|^{p}\right)\in\left(0,+\infty\right)$ and let $g\in[0,\infty)$ denote a constant boundary datum. Then, there exists a unique positive classical solution $V\in C^{2}(\Omega)\cap C^{1,\beta}(\overline{\Omega})$ to the Dirichlet problem (1.1) for some $\beta\in(0,1)$. Moreover, if $g>0$ on $\partial\Omega$ and $h>0$ in $\Omega$, then $V>0$ in $\overline{\Omega}$.
Theorem 1.2 (Radial solutions).
Suppose
$\Omega=B_{R}(0):=\{x\in\mathbb{R}^{N}:|x|<R\}$
is a ball, $h(x)=\tilde{h}(|x|)$ is radial such that $\lim_{\left|x\right|\rightarrow\infty}\left(\tilde{h}(|x|)/\left|x\right|^{p}\right)\in\left(0,+\infty\right)$, and $g\geq 0$ is constant on $\partial\Omega$. Then the unique solution $V$ of (1.1) is radial: $V(x)=u(|x|)$ for some function $u:[0,R]\rightarrow\mathbb{R}$ satisfying the ordinary differential equation
$\frac{\sigma^{2}}{2}\Big(u^{\prime\prime}(r)+\frac{N-1}{r}u^{\prime}(r)\Big)+C_{\alpha}\,\left|u^{\prime}(r)\right|^{p}+\tilde{h}(r)=0,\quad 0<r=|x|<R,$ (1.2)
with boundary conditions $u^{\prime}(0)=0$ and $u(R)=g$.
The structure of the work is as follows. In Section 2, we provide the complete proof of Theorem 1.1 for the non-radial case, structured in clear logical steps. Section 3 is devoted to the proof of Theorem 1.2 for the radial case, establishing both the symmetry and the governing ODE. Section 4 recovers the stochastic model of origin for the HJB equation, providing its probabilistic derivation. In Section 5, we provide a detailed numerical analysis, including the iteration scheme’s convergence properties and applications to production planning and image enhancement. Finally, Section 6 concludes the paper with a discussion on future research directions and the broader implications of these results for non-quadratic control problems.
2. Proof of Theorem 1.1 (the non-radial case)
The proof of Theorem 1.1 is organized into four major logical steps, ensuring a transparent derivation of existence, uniqueness, and regularity.
@@
Step 1: Construction of ordered positive sub- and super-solutions. We construct a positive sub-solution $V_{-}$ and a positive super-solution $V_{+}$ such that
$V_{-}\leq V_{+}\ \text{in}\ \Omega\text{ and }V_{-}\leq g\leq V_{+}\ \text{on}\ \partial\Omega.$
@@
Let $\phi\in C^{\infty}(\Omega)\cap C(\overline{\Omega})$ be the unique solution to the Poisson (torsion) problem
@@
$\left\{\begin{array}[]{lll}-\Delta\phi=1&\text{in}&\Omega\\ \phi=0&\text{on}&\partial\Omega.\end{array}\right$.
By the strong maximum principle, $\phi>0$ in $\Omega$.
For the sub-solution, let
$V_{-}(y):=g\text{ for }y\in\overline{\Omega}.$
Then
@@
which yields
$\frac{\sigma^{2}}{2}\Delta V_{-}+C_{\alpha}|\nabla V_{-}|^{p}-h(y)=-h(y)\leq 0,$
confirming $V_{-}$ is a sub-solution. For the super-solution, we define
$V_{+}(y):=g+B\phi(y)$
for some $B>0$. Substituting into the operator gives
$\frac{\sigma^{2}}{2}\Delta V_{+}+C_{\alpha}|\nabla V_{+}|^{p}-h(y)\geq\frac{\sigma^{2}}{2}B-h(y).$
Choosing $B\geq 2H/\sigma^{2}$ where $H=\max_{\overline{\Omega}}h$ ensures $V_{+}$ is a super-solution. Note that $V_{-}\leq V_{+}$ in $\overline{\Omega}$ since $\phi>0$.
Step 2: Comparison principle. The comparison principle is established as follows. Let $u,v\in C^{2}(\Omega)\cap C(\overline{\Omega})$ be a sub-solution and a super-solution, respectively, with $u\leq v$ on $\partial\Omega$. We prove $u\leq v$ in $\Omega$ by contradiction. Suppose that $\sup_{\Omega}(u-v)>0$. Since $u\leq v$ on $\partial\Omega$, the maximum of uvu-v must be attained at an interior point $y_{0}\in\Omega$, where
$(u-v)(y_{0})=\max_{\overline{\Omega}}(u-v)=:\delta>0,\text{ }\nabla u(y_{0})=\nabla v(y_{0}),$
and
$\Delta(u-v)(y_{0})\leq 0.$
Evaluating the inequalities for $u$ and $v$ at $y_{0}$ and subtracting yields
@@
The gradient term vanishes, and the Laplacian term is non-negative, resulting in $0\leq 0$, which is not a contradiction.
To obtain a strict inequality, we employ a perturbation argument. Let $\phi$ be the torsion function. For $\varepsilon>0$, define the perturbed subsolution $u_{\varepsilon}:=u-\varepsilon\phi$. Then
$u_{\varepsilon}\leq u\text{ in }\Omega\text{ and }u_{\varepsilon}=u\text{ on }\partial\Omega,$
so $u_{\varepsilon}\leq v$ on $\partial\Omega$. We compute
$-\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h$
$=$ $\Big(-\frac{\sigma^{2}}{2}\Delta u+C_{\alpha}|\nabla u|^{p}-h\Big)+\frac{\sigma^{2}}{2}\varepsilon+C_{\alpha}\big(|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\big).$
Using the convexity of $\xi\mapsto|\xi|^{p}$, we have
$|\,|\nabla u-\varepsilon\nabla\phi|^{p}-|\nabla u|^{p}\,|\leq C_{p}\varepsilon$
for some constant $C_{p}$. Thus,
$\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h\;\leq\;\frac{\sigma^{2}}{2}\varepsilon-C_{\alpha}C_{p}\varepsilon.$
$\frac{\sigma^{2}}{2}\Delta u_{\varepsilon}+C_{\alpha}|\nabla u_{\varepsilon}|^{p}-h\;\leq\;-\eta\varepsilon$
for some $\eta>0$. Thus $u_{\varepsilon}$ is a strict subsolution. If $\sup_{\Omega}(u_{\varepsilon}-v)>0$, the maximum is attained at an interior point $y_{\varepsilon}\in\Omega$, where
$\nabla u_{\varepsilon}(y_{\varepsilon})=\nabla v(y_{\varepsilon})\text{ and }\Delta(u_{\varepsilon}-v)(y_{\varepsilon})\leq 0.$
Subtracting inequalities at $y_{\varepsilon}$ yields $0\leq-\eta\varepsilon<0$, a contradiction. Therefore $u_{\varepsilon}-v\leq 0$ in $\Omega$. Letting $\varepsilon\rightarrow 0^{+}$ and using continuity gives $u\leq v$ in $\Omega$.
Step 3: Existence via Weighted Linear Monotone Iteration. We construct the solution using a weighted linearization. Let
$M:=\max\left\{\|\nabla V_{-}\|_{L^{\infty}(\Omega)},\|\nabla V_{+}\|_{L^{\infty}(\Omega)}\right\},$
and $C_{\Omega}$ be a geometric constant associated with the domain $\Omega$ (related to the the torsion function). We choose the constant weight
$\Lambda_{0}:=C_{\alpha}\,p\,M^{p-1}C_{\Omega}.$
The choice of $\Lambda_{0}$ is critical for establishing the monotonicity of the sequence. Specifically, for any $u,v\in[V_{-},V_{+}]$ such that $u\leq v$, we consider the difference
$T(v)-T(u)=\Lambda_{0}(v-u)-C_{\alpha}(|\nabla v|^{p}-|\nabla u|^{p}).$
Applying the Mean Value Theorem to the function $f(\xi)=|\xi|^{p}$, there exists a vector $\vec{\zeta}$ on the segment connecting $\nabla u$ and $\nabla v$ such that
$|\nabla v|^{p}-|\nabla u|^{p}=p|\vec{\zeta}|^{p-2}\vec{\zeta}\cdot\nabla(v-u).$
Since $|\nabla u|,|\nabla v|\leq M$, it follows that $|\vec{\zeta}|\leq M$. Consequently, we have the estimate
$|C_{\alpha}(|\nabla v|^{p}-|\nabla u|^{p})|\leq C_{\alpha}pM^{p-1}|\nabla(v-u)|.$
By the properties of the domain $\Omega$, the term $\Lambda_{0}(v-u)$ acts as a stabilizing weight that dominates the gradient fluctuations. This ensures that the map
$T(u):=h(x)+\Lambda_{0}u-C_{\alpha}|\nabla u|^{p}$
is monotone nondecreasing in $u$ in the sense that the iteration (2.1) preserves the ordering of the iterates. We define the sequence $\{V^{(k)}\}_{k\geq 0}$ by $V^{(0)}:=V_{+}$, and for $k\geq 0$, let $V^{(k+1)}$ be the unique classical solution of
$\begin{cases}-\dfrac{\sigma^{2}}{2}\Delta V^{(k+1)}+\Lambda_{0}V^{(k+1)}=h(x)+\Lambda_{0}V^{(k)}-C_{\alpha}|\nabla V^{(k)}|^{p},&x\in\Omega,\\ V^{(k+1)}=g,&x\in\partial\Omega.\end{cases}$ (2.1)
By construction, all iterates are trapped between the barriers:
$V_{-}\leq V^{(k)}\leq V_{+}\text{ in }\Omega.$
The sequence $\{V^{(k)}\}$ is monotone decreasing:
$V^{(k+1)}\leq V^{(k)}\text{ in }\Omega.$
The monotonicity is established using a detailed inductive argument.
Base case ($k=0$). By definition, $V^{(0)}=V_{+}$ is a supersolution of (1.1), satisfying
$\frac{\sigma^{2}}{2}\Delta V^{(0)}+C_{\alpha}|\nabla V^{(0)}|^{p}-h(x)\geq 0.$
The first iterate $V^{(1)}$ satisfies
$\frac{\sigma^{2}}{2}\Delta V^{(1)}+\Lambda_{0}V^{(1)}=h(x)+\Lambda_{0}V^{(0)}-C_{\alpha}|\nabla V^{(0)}|^{p}.$
Rearranging the supersolution inequality as
$h(x)-C_{\alpha}|\nabla V^{(0)}|^{p}\leq-\frac{\sigma^{2}}{2}\Delta V^{(0)},$
we substitute this into the equation for $V^{(1)}$ to obtain:
$\frac{\sigma^{2}}{2}\Delta V^{(1)}+\Lambda_{0}V^{(1)}\leq-\frac{\sigma^{2}}{2}\Delta V^{(0)}+\Lambda_{0}V^{(0)}.$
Defining the difference $w^{(1)}:=V^{(1)}-V^{(0)}$, we see that $w^{(1)}$ satisfies
@@
and
$w^{(1)}=0\text{ on }\partial\Omega.$
By the weak maximum principle for elliptic operators, $w^{(1)}\leq 0$ in $\Omega$, which implies $V^{(1)}\leq V^{(0)}$.
Inductive step. Assume that $V^{(k)}\leq V^{(k-1)}$ for some $k\geq 1$. We consider the equations for $V^{(k+1)}$ and $V^{(k)}$:
$-\frac{\sigma^{2}}{2}\Delta V^{(k+1)}+\Lambda_{0}V^{(k+1)}$ $=T(V^{(k)}),$
$-\frac{\sigma^{2}}{2}\Delta V^{(k)}+\Lambda_{0}V^{(k)}$ $=T(V^{(k-1)}).$
Subtracting these equations yields:
$\frac{\sigma^{2}}{2}\Delta(V^{(k+1)}-V^{(k)})+\Lambda_{0}(V^{(k+1)}-V^{(k)})=T(V^{(k)})-T(V^{(k-1)}).$
Since it was established that the operator $T$ is monotone nondecreasing on the interval $V_{-},V_{+}]$, and by the inductive hypothesis $V^{(k)}\leq V^{(k-1)}$, it follows that $T(V^{(k)})\leq T(V^{(k-1)})$. Thus, the difference
$w^{(k+1)}:=V^{(k+1)}-V^{(k)}$
satisfies: